# Not All Clients Are Equal: Personalized Federated Learning on Heterogeneous Multi-Modal Clients

### v1

@@ 6.1 Setups
Models.
To simulate model heterogeneity in federated MLLM training, we employ LLaVA-1.5 (liu2023visual, ) variants. Specifically, for LLM of LLaVA, we employ various sizes of Llama-3 (grattafiori2024llama, ) (i.e., Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B) and Qwen-2.5 (yang2024qwen2, ) (i.e., Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3B). For image encoder, we use CLIP ViT-L/336px (radford2021learning, ) in all variants, following  (radford2021learning, ; zhang2025llava, ).
Baselines.
We compare FedMosaic with state-of-the-art PFL methods. DITTO (li2021ditto, ), FedSim (pillutla2022federated, ), FedIT (zhang2024towards, ), TAKFL (morafah2024towards, ), FedDPA (long2024dual, ), FedDAT (chen2024feddat, ), PerAda (xie2024perada, ), and FedMKT (fan2025fedmkt, ). We also compare with supervised fine-tuning (SFT), which trains each client model independently without knowledge sharing and often performs better than participating in FL under data heterogeneity (ghari2024personalized, ).
Federated distillation baselines (i.e., TAKFL, PerAda, and FedMKT) share knowledge between heterogeneous models via logits, whereas FedMosaic employs PQ-LoRA. For the other baselines, we combine them with Fed-ET (cho2022heterogeneous, ), which aggregates models among homogeneous clients.
Metrics.
We report $A_{last}$, the accuracy at the end of training, and $A_{\text{AUC}}$ (koh2022online, ), which computes the area under the accuracy curve by measuring accuracy at each evaluation period to capture intermediate performance. For all experiments, we set the evaluation period to every five rounds and run them with 3 different random seeds, reporting the average performance along with the standard deviation.
Benchmarks.
We evaluate FedMosaic on our proposed DRAKE and HFLB (chen2024feddat, ), a multi-modal PFL benchmark. Our evaluation covers PFL-Static (i.e., i.i.d. client data distributions) and the more realistic PFL-Dynamic setup (i.e., intra-client distribution shifts with four incremental tasks). To evaluate the PFL-Dynamic setup on HFLB dataset, we partition the client data by question types (e.g., color, shape, count, and size). In addition, we evaluate on LLM-FL benchmarks, i.e., Fed-LLM (ye2024fedllm, ) and Fed-Aya (singh2024aya, ). See Sec. A.5 for details of the benchmarks and implementation details.
@@ 6.2 Quantitative Analysis
In all experiments, we evaluate each client’s model on its own task (i.e., ‘Self’) and on tasks from other clients (i.e., ‘Others’), reporting the average performance across clients. ‘Self’ performance shows personalization, while ‘Others’ indicates generalizability. Although personalization is the main goal of PFL, generalization ability is also crucial for continual personalization, as it enables rapid adaptation to newly arriving tasks in the future and addition of new clients during training (finn2017model, ; RAO2023271, ).
Method DRAKE-Dynamic HFLB-Dynamic
@@
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 65.79$\pm$0.20 57.62$\pm$0.26 47.66$\pm$0.13 46.67$\pm$0.11 80.65$\pm$0.41 77.79$\pm$0.08 61.49$\pm$0.09 61.17$\pm$0.13
DITTO (ICML 2021) 59.91$\pm$0.18 54.16$\pm$0.06 47.45$\pm$0.36 46.70$\pm$0.10 78.92$\pm$0.12 76.33$\pm$0.13 61.42$\pm$0.59 59.57$\pm$1.84
FedSim (ICML 2022) 63.98$\pm$1.19 56.42$\pm$0.70 47.04$\pm$0.16 46.15$\pm$0.17 79.30$\pm$0.26 76.60$\pm$0.25 59.48$\pm$0.80 58.84$\pm$0.81
FedIT (ICASSP 2024) 66.11$\pm$0.27 57.86$\pm$0.27 47.63$\pm$0.16 46.62$\pm$0.04 79.90$\pm$0.38 77.21$\pm$0.50 60.29$\pm$1.47 60.37$\pm$1.25
TAKFL (NeurIPS 2024) 64.54$\pm$0.85 56.19$\pm$0.98 47.38$\pm$0.06 46.47$\pm$0.14 79.23$\pm$0.35 76.75$\pm$0.46 61.04$\pm$0.62 59.62$\pm$1.91
FedDPA (NeurIPS 2024) 63.34$\pm$0.23 55.74$\pm$0.38 47.61$\pm$0.21 46.64$\pm$0.20 80.42$\pm$0.41 77.67$\pm$0.04 61.70$\pm$0.08 61.10$\pm$0.09
FedDAT (AAAI 2024) 58.47$\pm$1.10 54.38$\pm$0.28 48.91$\pm$0.42 47.78$\pm$0.18 78.98$\pm$0.38 76.45$\pm$0.62 64.82$\pm$0.54 64.16$\pm$0.39
PerAda (CVPR 2024) 59.75$\pm$1.06 54.48$\pm$0.13 47.30$\pm$0.45 46.72$\pm$0.03 78.61$\pm$0.66 75.92$\pm$0.42 61.24$\pm$0.95 60.83$\pm$0.06
FedMKT (COLING 2025) 61.38$\pm$0.18 55.45$\pm$0.31 47.50$\pm$0.07 46.68$\pm$0.08 78.70$\pm$0.38 75.94$\pm$0.68 61.12$\pm$0.26 59.59$\pm$1.86
FedMosaic (Ours) 67.86$\pm$0.51 59.83$\pm$0.16 51.16$\pm$0.04 49.36$\pm$0.08 80.90$\pm$0.30 79.18$\pm$1.04 66.66$\pm$0.77 65.80$\pm$0.30
Method DRAKE-Static HFLB-Static
@@
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 68.50$\pm$0.42 63.84$\pm$0.27 47.91$\pm$0.39 47.56$\pm$0.23 80.33$\pm$1.02 79.46$\pm$0.27 61.38$\pm$0.96 61.19$\pm$0.37
DITTO (ICML 2021) 63.67$\pm$1.50 59.17$\pm$1.25 48.22$\pm$0.09 47.65$\pm$0.03 79.65$\pm$0.87 78.23$\pm$0.25 61.98$\pm$0.37 61.38$\pm$0.08
FedSim (ICML 2022) 66.75$\pm$0.56 61.70$\pm$0.60 46.93$\pm$0.25 46.76$\pm$0.02 79.17$\pm$0.16 78.48$\pm$0.53 59.22$\pm$0.59 59.45$\pm$0.44
FedIT (ICASSP 2024) 68.71$\pm$0.04 63.89$\pm$0.55 47.91$\pm$0.22 47.60$\pm$0.17 80.08$\pm$1.07 79.41$\pm$0.17 61.32$\pm$0.75 61.16$\pm$0.27
TAKFL (NeurIPS 2024) 67.28$\pm$0.01 62.42$\pm$0.36 47.55$\pm$0.23 47.51$\pm$0.07 79.45$\pm$0.33 78.42$\pm$0.99 60.61$\pm$1.63 60.45$\pm$1.31
FedDPA (NeurIPS 2024) 66.09$\pm$1.51 61.36$\pm$0.14 47.93$\pm$0.30 47.69$\pm$0.11 79.60$\pm$1.37 78.55$\pm$0.50 61.33$\pm$0.25 60.94$\pm$0.01
FedDAT (AAAI 2024) 61.28$\pm$0.07 57.92$\pm$0.02 49.37$\pm$0.02 48.78$\pm$0.04 79.17$\pm$0.37 77.72$\pm$0.71 65.86$\pm$0.37 64.79$\pm$0.23
PerAda (CVPR 2024) 63.67$\pm$1.01 59.03$\pm$1.19 48.13$\pm$0.03 47.60$\pm$0.05 79.64$\pm$0.70 78.24$\pm$0.17 61.93$\pm$0.38 61.36$\pm$0.07
FedMKT (COLING 2025) 65.81$\pm$1.09 60.37$\pm$1.16 47.19$\pm$0.86 46.44$\pm$1.49 80.64$\pm$0.50 78.48$\pm$0.56 61.15$\pm$0.50 61.11$\pm$0.28
FedMosaic (Ours) 70.10$\pm$0.53 64.24$\pm$0.40 51.57$\pm$0.24 50.42$\pm$0.12 81.63$\pm$0.24 80.11$\pm$0.12 67.03$\pm$0.01 66.02$\pm$0.29
Table 2: Quantitative comparison in heterogeneous PFL. ‘Self’ denotes evaluation on a client’s own data, while ‘Others’ denotes evaluation on data from other clients. In DRAKE, 4 clients use LLaVA-Llama3.2-1B models and 6 clients use LLaVA-Llama3.2-3B models, while in HFLB, 3 clients use LLaVA-Llama3.2-1B models and 6 clients use LLaVA-Llama3.2-3B models. SFT refers to supervised fine-tuning on each client’s data without cross-client knowledge sharing.
Method $A_{last}$ / $A_{\text{AUC}}$
LLaVA-Llama3.2-1B LLaVA-Llama3.2-3B
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
FedMosaic (Ours) 77.42 / 60.45 58.50 / 55.26 66.71 / 59.60 63.13 / 55.84 68.87 / 63.60 77.61 / 69.80 69.19 / 56.66 70.54 / 58.93 70.68 / 54.09 59.21 / 55.32
Difference +0.97 / +0.57 +0.40 / +0.91 +0.08 / +0.04 +0.27 / +2.67 +6.86 / +2.09 +1.05 / +2.93 +6.61 / +2.03 +2.00 / -0.72 +4.76 / +2.16 +0.91 / +0.71
Table 3: Per-client performance of FedMosaic vs. SFT in a heterogeneous PFL-Dynamic setup on DRAKE. The Difference row shows the performance gain/loss of FedMosaic compared to SFT. Clients 1-4 use LLaVA-Llama3.2-1B, while Clients 5-10 use LLaVA-Llama3.2-3B.
Heterogeneous Multi-Modal Clients.
We first evaluate FedMosaic in multi-modal heterogeneous PFL-Dynamic and -Static setups, where clients employ different model architectures (LLaVA-Llama3-1B or 3B). As shown in Tab. 2, FedMosaic outperforms baselines on both clients’ own tasks (‘Self’) and others’ tasks (‘Others’) under both static and dynamic distributions. HFLB’s single VQA task with single-image inputs converges faster than DRAKE, leading to smaller gaps among baselines.
Note that the results in Tab. 2 report the average performance across all clients. Prior work has shown that local client training (i.e., SFT) may be sufficient for personalization in some tasks (wozniak2024personalized, ; mosbach2021on, ), which can make the overall average improvements appear small. For instance, as shown in the per-client accuracy in Tab. 3, FedMosaic shows only marginal improvements over SFT for clients 2 and 3. In contrast, for clients 5, 7, and 9, FedMosaic significantly improves personalization by sharing knowledge. We provide additional per-client results in Sec. A.12.
Moreover, as shown in Tab.3, not only do clients using smaller models (i.e., LLaVA-Llama3.2-1B) benefit from knowledge sharing through PFL, but larger models (i.e., LLaVA-Llama3.2-3B) also see significant gains. We attribute this to (i) SIMA accurately measuring task similarity under distribution shifts, reducing interference during aggregation, and (ii) PQ-LoRA enabling effective knowledge transfer between heterogeneous architectures, allowing smaller and larger models to mutually benefit.
We also emphasize that FedMosaic significantly outperforms baselines in generalization (‘Others’), which is crucial under distribution shifts where new, unseen tasks continuously emerge. This enhanced generalizability enables faster adaptation and accelerates future personalization, as shown in Fig. 5. The superior $A_{last}$ performance further confirms that FedMosaic effectively accumulates global knowledge during training, enabling quicker adaptation to new tasks.
Evaluation of Fast Adaptation Capability.
In real-world scenarios, unseen task data continuously emerge, making rapid adaptation crucial for future personalization. We evaluate adaptation speed by initializing models with aggregated local models from each PFL method and fine-tuning on unseen tasks for 200 iterations. As shown in Fig. 5, while sufficient training eventually leads to good downstream performance regardless of initialization, models initialized with FedMosaic achieve strong performance with just a few epochs, demonstrating that FedMosaic enhances generalizability by effectively sharing knowledge in heterogeneous PFL setups.
Heterogeneous PFL Setup in LLMs.
Beyond multi-modal PFL, we evaluate FedMosaic in heterogeneous PFL for LLM training. Tab. 4 shows FedMosaic significantly outperforms baselines in both ‘Self’ and ‘Others’, consistent with our MLLM-FL benchmark results.
Additional Experiments.
We compare accuracy and relative FLOPs in Fig. 6. As shown, FedMosaic demonstrates higher personalization performance with computation comparable to SFT.
We present experimental results for the Heterogeneous PFL setup using three heterogeneous architectures, LLaVA-Llama3.2-1B, LLaVA-Llama3.2-3B, and LLaVA-Llama3.1-8B (Sec.A.8), and LLaVA-Qwen2.5 variant models (Sec.A.10). We also present results in Homogeneous PFL setup (Sec. A.9), results in Fed-LLM benchmark (Sec. A.11), extended fast adaptation evaluations (Sec. A.17), computation analysis (Sec. A.6), details of client model selection in the Heterogeneous PFL setup(Sec. A.3).
Method Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
SFT 2.34$\pm$0.08 2.17$\pm$0.03 2.01$\pm$0.02 1.99$\pm$0.01
FedSim (ICML 2022) 1.99$\pm$0.03 1.98$\pm$0.04 1.92$\pm$0.01 1.92$\pm$0.03
FedIT (ICASSP 2024) 2.27$\pm$0.08 2.18$\pm$0.02 2.19$\pm$0.11 2.10$\pm$0.04
TAKFL (NeurIPS 2024) 2.34$\pm$0.12 2.17$\pm$0.06 2.19$\pm$0.10 2.08$\pm$0.03
FedDPA (NeurIPS 2024) 2.39$\pm$0.09 2.27$\pm$0.02 2.20$\pm$0.05 2.13$\pm$0.01
PerAda (CVPR 2024) 2.33$\pm$0.09 2.12$\pm$0.03 2.12$\pm$0.10 1.98$\pm$0.02
FedMKT (COLING 2025) 2.31$\pm$0.06 2.12$\pm$0.05 2.06$\pm$0.02 1.96$\pm$0.00
FedMosaic (Ours) 2.51$\pm$0.01 2.32$\pm$0.03 2.25$\pm$0.01 2.17$\pm$0.02
Table 4: Quantitative comparison in heterogeneous LLM Clients on FedAya-Dynamic. We compute GPT scores evaluated by GPT-4 (Sec. A.5).
Figure 6: Accuracy and relative FLOPs in heterogeneous DRAKE-Dynamic.
6.3 Ablation Study
We ablate FedMosaic to investigate the benefit of each proposed component, and summarize the results in Tab. 5. Our observations indicate that each component plays a significant role in enhancing both the personalization (‘Self’) and the generalization (‘Others’) performance. Specifically, PQ-LoRA enhances generalization by transferring knowledge across heterogeneous architectures, while SIMA improves personalization by promoting knowledge sharing among similar-task models only.
We further present an ablation study: effect of low rank $r$ in PQ-LoRA (Sec. A.13), effect of $N_{B}$ (Sec. A.16), effect of weight alignment in PQ-LoRA (Sec. A.15), and the effect of decayed client-wise gradient $\hat{g}$ in SIMA (Sec. A.14). We discuss limitations in Sec. A.19.
Method DRAKE HFLB
Self Others Self Others
$A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{last}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$
Vanilla 66.10$\pm$0.27 57.87$\pm$0.27 47.63$\pm$0.16 46.62$\pm$0.04 79.90$\pm$0.38 77.21$\pm$0.50 60.29$\pm$1.47 60.37$\pm$1.25
(+) PQ-LoRA 66.99$\pm$0.77 58.39$\pm$0.38 51.33$\pm$0.06 48.78$\pm$0.03 80.55$\pm$0.11 77.86$\pm$0.45 66.63$\pm$0.30 65.92$\pm$0.08
(+) PQ-LoRA &SIMA (Ours) 67.86$\pm$0.51 59.83$\pm$0.16 51.16$\pm$0.04 49.36$\pm$0.08 80.90$\pm$0.30 79.18$\pm$1.04 66.66$\pm$0.77 65.80$\pm$0.30

### v3

@@ 6.1 Setups
Models. To simulate model heterogeneity in federated MLLM training, we employ LLaVA-1.5 (Liu et al., 2023b) variants. For LLM of LLaVA, we employ various sizes of Llama-3 (Grattafiori et al., 2024) (Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B) and Qwen-2.5 (Yang et al., 2024a) (Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3B). We use the Llama-3 series for text-only benchmarks.
Metrics. We report $A_{last}$, the accuracy at the end of training, and $A_{\text{AUC}}$ (Koh et al., 2022), which computes the area under the accuracy curve by measuring accuracy at each evaluation period to capture intermediate performance. All experiments use five rounds of evaluation intervals and are averaged over three different random seeds, with standard deviations reported.
Benchmarks. We evaluate FedMosaic on multi-modal PFL benchmarks: HFLB (Chen et al., 2024) and our proposed DRAKE. Our evaluation covers PFL-Static (i.e., i.i.d. client data distributions) and the more realistic PFL-Dynamic setup (i.e., intra-client distribution shifts with four incremental tasks). For PFL-Dynamic on HFLB, we partition client data by question types (e.g., color, shape, count, and size). We also evaluate on text-only PFL benchmarks, i.e., Fed-Scope (Kuang et al., 2024), Fed-Aya (Singh et al., 2024), and Fed-LLM-Large that we combined Fed-LLM (Ye et al., 2024) and Fed-FLAN (Long et al., 2024). See Sec. A.6 for the details of experiment setup.
Baselines. We compare FedMosaic with SOTA PFL methods: DITTO (Li et al., 2021), FedSim (Pillutla et al., 2022), FedIT (Zhang et al., 2024a), TAKFL (Morafah et al., 2024), FedDPA (Long et al., 2024), FedDAT (Chen et al., 2024), PerAda (Xie et al., 2024), and FedMKT (Fan et al., 2025). We also compare with supervised fine-tuning (SFT), where clients train independently without knowledge sharing and often outperforming FL methods under data heterogeneity (Ghari & Shen, 2024).
@@ 6.2 Quantitative Analysis
In all experiments, we evaluate each client’s model on its own task (i.e., ‘Self’) and on other clinets’ tasks (i.e., ‘Others’), reporting average performance across clients. ‘Self’ performance shows personalization, while ‘Others’ indicates generalizability. Although personalization is PFL’s main goal, generalization ability is also crucial for continual personalization, as it enables rapid adaptation to new tasks in the future during training (Finn et al., 2017; Rao et al., 2023).
Heterogeneous Multi-Modal Clients. We evaluate FedMosaic in multi-modal heterogeneous PFL-Dynamic and -Static setups, where clients employ different architectures (LLaVA-Llama3-1B or 3B). As shown in Tab. 2 and Tab. 4, FedMosaic consistently outperforms baselines on both clients’ own tasks (‘Self’) and others’ tasks (‘Others’) under static and dynamic distributions. HFLB’s single VQA task with single-image inputs converges faster than DRAKE, yielding smaller gaps among baselines.
DRAKE-Dynamic HFLB-Dynamic
@@
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 2: Quantitative comparison in heterogeneous PFL. ‘Self’ denotes evaluation on a client’s own data, while ‘Others’ denotes evaluation on data from other clients. In DRAKE, 4 clients use LLaVA-Llama3.2-1B and 6 clients use LLaVA-Llama3.2-3B, while in HFLB, 3 clients use LLaVA-Llama3.2-1B and 6 clients use LLaVA-Llama3.2-3B. SFT refers to supervised fine-tuning on each client’s data without cross-client knowledge sharing.
Self Others
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 3: Quantitative comparison in PFL on DRAKE-static. 4 clients use LLaVA-Llama3.2-1B, 6 clients use LLaVA-Llama3.2-3B.
Self Others
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 4: Quantitative comparison in cross-family PFL on DRAKE-dynamic. 3 clients use LLaVA-Llama3.2-3B, 4 use LLaVA-Qwen2.5-1.5B, 3 use LLaVA-Qwen2.5-3B.
Self $_{last} / AA$$_{\text{AUC}}$
LLaVA-Llama3.2-1B LLaVA-Llama3.2-3B
Method Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 Client 7 Client 8 Client 9 Client 10
FedMosaic 77.42 / 60.45 58.50 / 55.26 66.71 / 59.60 63.13 / 55.84 68.87 / 63.60 77.61 / 69.80 69.19 / 56.66 70.54 / 58.93 70.68 / 54.09 59.21 / 55.32
Gain +0.97 / +0.57 +0.40 / +0.91 +0.08 / +0.04 +0.27 / +2.67 +6.86 / +2.09 +1.05 / +2.93 +6.61 / +2.03 +2.00 / -0.72 +4.76 / +2.16 +0.91 / +0.71
Table 5: Per-client ‘Self’ performance of FedMosaic vs. SFT in a heterogeneous PFL-Dynamic setup on DRAKE. The Gain row shows positive gain of FedMosaic in blue, negative in red against SFT. Clients 1-4 use LLaVA-Llama3.2-1B, while Clients 5-10 use LLaVA-Llama3.2-3B.
Note that Tab. 2 reports the average performance across all clients. Local client training (SFT) can suffice for simpler tasks (Mosbach et al., 2021; Woźniak et al., 2024), e.g., single-image VQA (clients 2, 3 in Tab. 5), making average improvements appear small. In contrast, FedMosaic significantly enhances personalization for complex multi-image tasks (clients 5, 7, 9 in Tab. 5) through effective knowledge sharing. See Sec.A.16 for additional per-client results and Sec. A.24 for task descriptions.
As shown in Tab.5, not only do clients using smaller models (i.e., LLaVA-Llama3.2-1B) benefit from knowledge sharing through PFL, but larger models (i.e., LLaVA-Llama3.2-3B) also see significant gains. We attribute this to (i) RELA accurately measuring task relevance under distribution shifts, reducing interference during aggregation, and (ii) PQ-LoRA enabling effective knowledge transfer between heterogeneous architectures, allowing smaller and larger models to mutually benefit.
We also emphasize that FedMosaic significantly outperforms baselines in generalization (‘Others’), which is crucial under distribution shifts where new, unseen tasks continuously emerge. This enhanced generalizability enables faster adaptation and accelerates future personalization, as shown in Fig. 6.
Cross-family Heterogeneity. Beyond varying model sizes within the same family (e.g., LLaVA-Llama-1B/3B), we further simulate cross-family heterogeneity (e.g., Qwen- vs. Llama-based MLLMs). Tab. 4 shows FedMosaic consistently outperforms baselines under cross-family heterogeneity, highlighting its generality. See Sec. A.14 for additional cross-family experiments.
Fast Adaptation Evaluation. In real-world scenarios, unseen tasks continuously emerge, making rapid adaptation crucial for future personalization. We evaluate adaptation speed by initializing models with each PFL method’s aggregated model and fine-tuning on unseen tasks for 200 iterations. Fig. 6 shows that while sufficient training eventually yields similarly high performance regardless of initialization, models initialized with FedMosaic achieve high performance within few steps, highlighting enhanced generalizability through effective knowledge sharing in heterogeneous PFL.
Large-Scale Evaluation of Heterogeneous PFL in LLMs. Beyond multi-modal PFL, we evaluate FedMosaic in heterogeneous PFL for LLM training (i.e., text-only NLP domain). Tab. 6 shows FedMosaic significantly outperforms baselines in both ‘Self’ and ‘Others’, consistent with our MLLM-FL benchmark results. This experiment involves 52 clients, highlighting the scalability of FedMosaic to large client populations and its applicability in real-world deployments.
Computation and Communication Cost Analysis. We compare accuracy and relative FLOPs in Fig. 7. FedMosaic demonstrates higher personalization performance with computation comparable to SFT. See Sec. A.8 and Sec. A.9 for detailed analyses of computation and communication costs.
Additional Experiments. Additional results include: heterogeneous PFL with LLaVA-Llama3.2-1B/3B/8B (Sec.A.11), LLaVA-Qwen2.5 variants (Sec.A.13), text-only benchmarks (Sec. A.15), homogeneous PFL setup (Sec. A.12), extended fast adaptation evaluations (Sec. A.23), comparison of RELA with similarity-aware aggregation methods (Sec. A.22), and client model selection details (Sec. A.3) and configurations (Sec. A.7). We further provide hyperparameter analysis Sec. A.18), effect of weight alignment in PQ-LoRA (Sec. A.21), effect of client model assignment (Sec. A.4), effect of decayed client-wise gradient $\hat{g}$ in RELA (Sec. A.19), and limitations (Sec. A.26).
Self Others
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]
Table 6: Large-Scale Experiments with 52 heterogeneous LLMs (Llama-1B/3B) on Fed-LLM-Large.
Figure 7: Accuracy and relative FLOPs in DRAKE-Dynamic.
Ablation Study. We ablate FedMosaic to investigate each proposed component’s benefit, and summarize the results in Tab. 7. Our observations indicate that each component significantly enhances both the personalization (‘Self’) and the generalization (‘Others’) performance. Specifically, PQ-LoRA enhances generalization by transferring knowledge across heterogeneous architectures, while RELA improves personalization by promoting knowledge sharing among relevant-task models only.
DRAKE Fed-Scope
@@
Method $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$ $_{last}\ \uparrow AA$$_{\text{AUC}}\ \uparrow$
[…]